4. Find the square roots of the following numbers by the Prime Factorisation Method.  v. 7744  

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student



7744 = 2×2×2×2×2×2×11×11×1
⇒ 7744 = (2×2)×(2×2)×(2×2)×(11×11)
⇒ 7744 = (2×2×2×11)×(2×2×2×11)
⇒ 7744 = (2×2×2×11)2
⇒ √7744 = 2×2×2×11 = 88